# LiteGE: Lightweight Geodesic Embedding for Efficient Geodesics Computation and Non-Isometric Shape Correspondence

### v1

@@ Abstract
Computing geodesic distances on 3D surfaces is fundamental to many tasks in 3D vision and geometry processing, with deep connections to tasks such as shape correspondence. Recent learning-based methods achieve strong performance but rely on large 3D backbones, leading to high memory usage and latency, which limit their use in interactive or resource-constrained settings. We introduce LiteGE, a lightweight approach that constructs compact, category-aware shape descriptors by applying PCA to unsigned distance field (UDFs) samples at informative voxels. This descriptor is efficient to compute and removes the need for high-capacity networks. LiteGE remains robust on sparse point clouds, supporting inputs with as few as 300 points, where prior methods fail. Extensive experiments show that LiteGE reduces memory usage and inference time by up to 300 compared to existing neural approaches. In addition, by exploiting the intrinsic relationship between geodesic distance and shape correspondence, LiteGE enables fast and accurate shape matching. Our method achieves up to 1000 speedup over state-of-the-art mesh-based approaches while maintaining comparable accuracy on non-isometric shape pairs, including evaluations on point-cloud inputs.
@@ Introduction
Due to the efficiency of our compact UDF representation, LiteGE supports rapid inference of correspondences for individual query points, making it particularly suitable for local region matching or sparse landmark identification on high-resolution meshes. Experiments conducted on the SMAL dataset of 4-legged animals demonstrate that LiteGE achieves an accuracy comparable to the state-of-the-art Spectral-Meets-Spatial (SMS) (cao2024spectral), which is a mesh-based method, while providing inference speeds that are 100 to 1000 faster. Notably, thanks to our novel shapes canonicalization pipeline, we can generalize the efficiency and accuracy of our method from meshes to point clouds.
@@
As shown in the figure, the mesh-based normalization used in shape matching achieves slightly higher explained variance than the point cloud-based normalization for geodesic regression. This is expected since mesh-based normalization employs unit-area scaling and TNet alignment, which are more consistent than PCA axis alignment and bounding-box scaling for point clouds. PCA alignment is sensitive to axis flipping, and bounding-box scaling is less precise. This supports the observation that better normalization reduces data variability and effective dimensionality.

### v2

@@ Abstract
Computing geodesic distances on 3D surfaces is fundamental to many tasks in 3D vision and geometry processing, with deep connections to tasks such as shape correspondence. Recent learning-based methods achieve strong performance but rely on large 3D backbones, leading to high memory usage and latency, which limit their use in interactive or resource-constrained settings. We introduce LiteGE, a lightweight approach that constructs compact, category-aware shape descriptors by applying Principal Component Analysis (PCA) to Unsigned Distance Field (UDFs) samples at informative voxels. This descriptor is efficient to compute and removes the need for high-capacity networks. LiteGE remains robust on sparse point clouds, supporting inputs with as few as 300 points, where prior methods fail. Extensive experiments show that LiteGE reduces memory usage and inference time by up to 300 compared to existing neural approaches. In addition, by exploiting the intrinsic relationship between geodesic distance and shape correspondence, LiteGE enables fast and accurate shape matching. Our method achieves up to 1000 speedup over state-of-the-art mesh-based approaches while maintaining comparable accuracy on non-isometric shape pairs, including evaluations on point-cloud inputs.
@@ Introduction
Due to the efficiency of our compact UDF representation, LiteGE supports rapid inference of correspondences for individual query points, making it particularly suitable for local region matching or sparse landmark identification on high-resolution meshes. Experiments conducted on the Skinned Multi-Animal Linear Model (SMAL) dataset of 4-legged animals demonstrate that LiteGE achieves an accuracy comparable to the state-of-the-art Spectral-Meets-Spatial (SMS) (cao2024spectral), which is a mesh-based method, while providing inference speeds that are 100 to 1000 faster. Notably, thanks to our novel shapes canonicalization pipeline, we can generalize the efficiency and accuracy of our method from meshes to point clouds.
@@
As shown in the figure, the mesh-based normalization used in shape matching achieves slightly higher explained variance than the point cloud-based normalization for geodesic regression. This is expected since mesh-based normalization employs unit-area scaling and T-Net alignment, which are more consistent than PCA axis alignment and bounding-box scaling for point clouds. PCA alignment is sensitive to axis flipping, and bounding-box scaling is less precise. This supports the observation that better normalization reduces data variability and effective dimensionality.